# Melodia: Training-Free Music Editing Guided by Attention Probing in Diffusion Models

### v1

@@ Abstract
Codehttps://aaai.org/example/code
Extended versionhttps://aaai.org/extended-version
@@ INTRODUCTION
To address these issues, we conduct an in-depth probing analysis on different attention maps within AudioLDM 2 (Liu et al. 2024c), a diffusion-based model commonly used as the backbone for existing music editing methods. We reveal a key insight: cross-attention maps encompass details regarding distinct musical characteristics, and interventions on these maps result in ineffective modifications. In contrast, self-attention maps are essential for preserving the temporal structure of the source music during its conversion into the target music. This finding suggests that manipulating self-attention maps may yield better editing results. Building on this insight, we introduce Melodia, a training-free and source-prompt-free approach that selectively manipulates self-attention maps in specific layers during the diffusion model’s denoising process. By building an attention repository to store self-attention information from the source music and applying it during editing, our method achieves a good balance between textual adherence and structural integrity without requiring any textual description of the source music. Fig. 1 demonstrates that our proposed method Melodia achieves superior structural preservation compared to existing approaches. Furthermore, to address the challenge of evaluating the balance between textual adherence and structural integrity, we propose two novel metrics, Adherence-Structure Balance Score (ASB) and Adherence-Musicality Balance Score (AMB) that comprehensively measure this balance in music editing tasks.
@@
  • We design two new evaluation metrics, ASB and AMB, to assess the balance between adherence to the target prompt and structural integrity in music editing tasks.
@@
AudioLDM 2 (Liu et al. 2024c), an open-source conditional LDM-based audio generation model, is the foundation model of our work. Detailed introduction in the Appendix.
@@
Probing Methodology. In AudioLDM 2 (Liu et al. 2024c), attention mechanisms are the key to its successful music generation. However, it remains unclear whether these attention maps are merely weight matrices or contain rich semantic representations of music. To address this, inspired by probing analysis methods in NLP (Clark et al. 2019; Liu et al. 2019), we build datasets and train classification networks to explore attention map properties. The principle is straightforward: If a classifier can accurately categorize attention maps into different classes, these maps must encode meaningful semantic information beyond mere weighting. We construct three prompt datasets targeting fundamental musical dimensions: instruments (16 types, a solo [instrument] music), styles (11 types, a typical [style] music), and moods (8 types, a [mood] music). For each prompt, we employ probing methods to extract cross-attention maps corresponding to specific keywords ([instrument], [style], and [mood]) and self-attention maps from different layers during generation. Finally, we train a simple MLP classifier to determine if these maps encode specific musical attributes. High classification accuracy indicates that attention maps encode substantial information about musical attributes rather than functioning merely as weighting mechanisms. For details about our probing method and prompt datasets, please refer to Appendix.
@@
Probing Results on Cross-Attention Maps. Tab. 1 shows our classification results on cross-attention maps across different layers, displaying partial results due to space constraints. The table demonstrates that our classifier achieves high accuracy across all three classification tasks, with accuracy rates exceeding 70% for all categories shown, indicating that cross-attention maps contain rich semantic information about musical characteristics rather than simple weighting. This explains why direct manipulation of cross-attention maps may lead to unsuccessful music editing outcomes. Fig. 2 shows that when replacing cross-attention maps (layers 8-14 or 1-16), the temporal structure of the edited music significantly differs from the source music, yet aligns closely with the music directly generated from the target prompt. For complete experimental results, please refer to the Appendix.
Probing Results on Self-Attention Maps. Tab. 2 shows that our classifier struggles with self-attention maps, achieving low accuracy (<<40%) significantly below cross-attention results. This suggests self-attention maps do not encode categorical information about musical attributes. Instead, our further analysis reveals they capture temporal features critical for music coherence, such as melody and rhythm. This finding parallels observations in image diffusion models, where self-attention maps preserve the spatial structure of images (Liu et al. 2024a). To validate this, we conducted self-attention map replacement experiments. Fig. 2 shows that replacing self-attention maps in layers 8-14 successfully changes attributes while preserving original melody and rhythm—the edited violin music maintains the source drum’s fast beat pattern. Without replacement, the music loses source temporal structure and resembles direct generation from the target prompt. However, replacing all layers (1-16) partially preserves original timbre instead of complete transformation to violin. These experimental results support our idea that self-attention maps are essential for preserving temporal structure during music editing, with selective layer replacement showing promising results.
@@
Based on our exploration of attention layers, the fundamental innovation of our approach lies in preserving the temporal structure of the original music clip $x_{0}$ within self-attention map manipulation. Based on this, we propose a straightforward yet effective approach named Melodia.
As shown in Fig. 3, in our approach, we first obtain the latent $z_{0}$ of the original music clip via a VAE (Kingma, Welling et al. 2013) encoder $\mathcal{E}$, and then collect the SA features of $z_{0}$ over the Partial DDIM Inversion (Song, Meng, and Ermon 2020; Manor and Michaeli 2024) process (for the sake of brevity, Partial DDIM inversion will be referred to as DDIM inversion in subsequent sections), a version that applies the inversion only up to $T_{\rm start}\in[0,T],T=1000$ (Manor and Michaeli 2024). Specifically, for predefined time steps $t\in\{0,\dots,T_{\rm start}\}$, we invert the original latent $z_{0}$ to an editable intermediate latent $z_{T_{\rm start}}$. And we store self-attention queries $Q^{\rm s}_{t}$ and keys $K^{\rm s}_{t}$ at each time step to build an attention repository in this process. The repository is used to provide guidance for subsequent music editing. In the music editing process, we start by transforming the information stored in the attention repository into guidance through the proposed Attention-based Structure Retention (ASR) method, which is built based on manipulation of the self-attention mechanisms of pre-trained diffusion models, without extra fine-tuning or new blocks. Similar to the guidance of the given target prompt, this guidance will also modulate the denoising process. Finally, the output $z^{\prime}_{0}$ will be decoded to a readable mel spectrogram of the desired music clip via the respective VAE decoder $\mathcal{D}$. See the pseudocode of our algorithm in the Appendix.
@@
Based on the above assumption, an intuitive illustration is presented in Fig. 4, showing a Structure-Timbre sampling space of the music latent diffusion model (Liu et al. 2023, 2024c). In our music editing process, the starting latent $z^{\prime}_{T_{\rm start}}$, copied from DDIM Inversion (Song, Meng, and Ermon 2020) prior $z_{T_{\rm start}}$, provides implicit structure guidance for generation. However, as shown in Fig. 4 (a), relying solely on this implicit guidance leads to significant structure divergence in the editing process (Song, Meng, and Ermon 2020; Mou et al. 2023; Tumanyan et al. 2023). Since the target prompt $y$ contains rich semantics of music, it can provide strong semantic guidance for editing, which is significantly stronger than the implicit structure guidance. Therefore, we build an attention repository to store SA features at each inversion time step in the system memory, and use the stored features to provide explicit structure guidance for the editing process, minimizing structure divergence (shown in Fig. 4 (b)). See details of GPU memory and inference time overhead of attention repository in the Appendix.
@@
where $\phi^{\prime\rm s}_{t}$ is the output of a manipulated self-attention layer and $V^{\prime\rm s}_{t}$ is the projected self-attention values of target latent $z^{\prime}_{t}$ at the specific time step $t$. In addition, we apply this manipulation to layers 8-14 of AudioLDM 2. The analysis of this layer selection strategy is provided in the Experiments section.
@@
MusicDelta zoME-Bench Supplementary Dataset
@@
For subjective evaluation, we introduce the Music Editing Balance (MEB) metric to capture human perception. Additionally, adapted from MusicMagus (Zhang et al. 2024b), we employ Relevance (REL) to assess how well the edited music aligns with the target prompt, and Structural Consistency (CON) to evaluate the consistency of the pitch contour and structural aspects. These metrics are rated on a five-point Likert scale (Likert 1932), with higher scores being better. More details of these metrics can be found in Appendix.
Figure 5: Quantitative Comparison with methods over $T_{\text{start}}$ range of 300-1000 on supplementary dataset. The highlighted region is the optimal balance region where shows both text adherence and structural integrity. Our method outperforms other approaches across all $T_{\text{start}}$ values.
@@
We apply Z-score normalization $\mathcal{N}(\cdot)$ and Min-Max scaling $\mathcal{S}(\cdot)$ to ensure scores fall within $[0,1]$, with 1 indicating best performance and 0 the worst across evaluated methods. Detailed procedures are provided in the Appendix.
@@
We evaluate our method using two existing datasets and a self-constructed supplementary dataset. We use the MusicDelta subset from the MedleyDB dataset (Bittner et al. 2014) and adopt the corresponding prompt dataset MedleyMDPrompts created by Manor and Michaeli (2024). From the ZoME-Bench dataset (Liu et al. 2024b), we select three subsets targeting instrument, genre, and mood changes for intra-stem editing tasks. We also construct a comprehensive supplementary dataset comprising both synthesized and authentic music recordings, with multiple subsets designed for the same three editing tasks. Detailed dataset specifications are provided in the Appendix.
@@
We evaluate our proposed method through a comparison with five competing approaches. All diffusion-based methods are implemented using the pre-trained AudioLDM 2 (Liu et al. 2024c) model with 200 inference steps. The target CFG of our method is set to 5.5. Details on CFG analysis and experimental settings can be found in Appendix.
Quantitative Comparison with baselines on datasets. Table 3 shows the performance comparison of our method with baseline approaches. Our method achieves competitive CLAP and the lowest LPAPS across all datasets, indicating superior text adherence and temporal structure preservation. Melodia also maintains good Chroma and FAD performance, effectively preserving harmonic elements and audio quality. While MusicMagus (Zhang et al. 2024b) shows high Chroma and FAD on the supplementary dataset, this indicates editing failure as outputs remain nearly unchanged from the source music, explaining its poor textual adherence (CLAP: 0.27). The ASB and AMB results demonstrate the necessity of our proposed composite metrics. Several baselines achieve 0.00 on these metrics due to severe imbalances between text adherence and structure preservation, while Melodia achieves excellent performance on both composite metrics, demonstrating superior balance between editing effectiveness and structure preservation. Task-specific results within each dataset are provided in the Appendix.
@@
MusicDelta zoME-Bench Supplementary
@@
Quantitative Comparison with baselines across $T_{\text{start}}$s. To quantitatively evaluate the adherence of edited music to the target prompt and the fidelity to the original music, we plot the CLAP-LPAPS results on our supplementary dataset for all methods in Fig. 5. MusicMagus (Zhang et al. 2024b) is tested only at $T_{\text{start}}=1000$ following its original setting, while other diffusion-based methods are evaluated across multiple $T_{\text{start}}$ values ranging from 300-1000. The results show that MusicGen (Copet et al. 2023) achieves good textual adherence but poor structural preservation. Notably, other diffusion-based methods also fail to precisely preserve the structure of the source music. In contrast, Melodia addresses this issue by introducing structure guidance and achieves better LPAPS scores for any level of CLAP.
@@
Qualitative Evaluation. Fig. 6 shows spectrogram comparisons of editing results between Melodia and baseline methods. Our method successfully preserves the temporal structure of the original music while effectively transferring the target attributes, which can be verified by more consistent rhythmic patterns maintained from the source music and successful frequency pattern transfer. For fair comparison, all methods use the same hyperparameter settings as in the quantitative evaluation.
@@
We implemented subjective listening evaluations across three datasets. A total of 100 participants were recruited from the MIR community and music forums, with 77 valid responses retained after screening. All participants had some degree of musical experience. The test is approved by the Institutional Review Board (IRB). We randomly selected 10 test samples from each dataset, and participants rated the edited results from each method. As shown in Table 4, Melodia substantially outperforms all baseline methods on the MEB metric, indicating that our method successfully achieves the desired balance between textual alignment and structural preservation. Melodia also obtained high scores on REL and CON metrics, confirming its superior performance in semantic alignment and structure preservation. For more details, please refer to the Appendix.
@@
Effects of layer selection. The selection of SA replacement layers is crucial in balancing structure preservation and attribute modification. Experimental results on the supplementary dataset demonstrate that replacing SA maps in layers 8–14 of AudioLDM 2 achieves optimal performance. As shown in Tab. 5, this layer selection yields the highest balance scores (ASB and AMB) while maintaining low structural disruption (LPAPS) and high textual alignment (CLAP). Visual evidence in Fig. 2 also confirms that SA manipulation in layers 8-14 successfully preserves rhythmic patterns and melodic contours during attribute transfer.
Results on another pre-trained diffusion model. To evaluate the generalization capability of Melodia across different diffusion models, we conducted experiments with the Stable Audio Open model (Evans et al. 2025) for timbre transfer on our supplementary dataset. As shown in Tab. 6, integrating Melodia with Stable Audio Open achieves improvements in both textual adherence (CLAP) and structural integrity (LPAPS and Chroma) compared to the baseline. These results confirm Melodia’s generalizability across different model architectures.

### v2

@@ INTRODUCTION
To address these issues, we conduct an in-depth probing analysis on different attention maps within AudioLDM 2 (Liu et al. 2024c), a diffusion-based model commonly used as the backbone for existing music editing methods. We reveal a key insight: cross-attention maps encompass details regarding distinct musical characteristics, and interventions on these maps result in ineffective modifications. In contrast, self-attention maps are essential for preserving the temporal structure of the source music during its conversion into the target music. This finding suggests that manipulating self-attention maps may yield better editing results. Building on this insight, we introduce Melodia, a training-free and source-prompt-free approach that selectively manipulates self-attention maps in specific layers during the diffusion model’s denoising process. The source-prompt-free nature lowers the usage barrier for non-expert users. By building an attention repository to store self-attention information from the source music and applying it during editing, our method achieves a good balance between textual adherence and structural integrity without requiring any textual description of the source music. Fig. 1 demonstrates that our proposed method Melodia achieves superior structural preservation compared to existing approaches. Furthermore, to address the challenge of evaluating the balance between textual adherence and structural integrity, we propose two novel metrics, Adherence-Structure Balance Score (ASB) and Adherence-Musicality Balance Score (AMB) that comprehensively measure this balance in music editing tasks. Additionally, we construct MelodiaEdit, a music editing benchmark covering diverse editing scenarios.
@@
  • We propose two novel evaluation metrics, ASB and AMB, to assess the balance between textual adherence and structural integrity, and construct MelodiaEdit benchmark covering diverse editing scenarios.
@@
AudioLDM 2 (Liu et al. 2024c), an open-source conditional LDM-based audio generation model, is the foundation model of our work.It contains 16 layers with cross-attention and self-attention mechanisms in its UNet architecture.
@@
Probing Methodology. In AudioLDM 2 (Liu et al. 2024c), attention mechanisms are the key to its successful music generation. However, it remains unclear whether these attention maps are merely weight matrices or contain rich semantic representations of music. To address this, inspired by probing analysis methods in NLP (Clark et al. 2019; Liu et al. 2019), we build datasets and train classification networks to explore attention map properties. The principle is straightforward: If a classifier can accurately categorize attention maps into different classes, these maps must encode meaningful semantic information beyond mere weighting. We construct three prompt datasets targeting fundamental musical dimensions: instruments (16 types, a solo [instrument] music), styles (11 types, a typical [style] music), and moods (8 types, a [mood] music). For each prompt, we extract cross-attention maps corresponding to specific keywords ([instrument], [style], and [mood]) and self-attention maps from different layers during generation. Finally, we train a simple MLP classifier to determine if these maps encode specific musical attributes. High classification accuracy indicates that attention maps encode substantial information about musical attributes rather than functioning merely as weighting mechanisms.
@@
Probing Results on Cross-Attention Maps. Tab. 1 shows our classification results on cross-attention maps across different layers. Our classifier achieves high accuracy across all three classification tasks, with average accuracy rates exceeding 70%, indicating that cross-attention maps contain rich semantic information about musical characteristics rather than simple weighting. This explains why direct manipulation of cross-attention maps causes unsuccessful music editing outcomes. Fig. 2 shows that when replacing cross-attention maps (layers 8-14 or 1-16), the temporal structure of the edited music significantly differs from the source music, yet aligns closely with the music directly generated from the target prompt.
Probing Results on Self-Attention Maps. Tab. 2 shows that our classifier struggles with self-attention maps, achieving low average accuracy (<<40%) significantly below cross-attention results. This suggests self-attention maps do not encode categorical information about musical attributes. Instead, our further analysis shows they capture temporal features critical for music coherence, such as melody and rhythm. This finding parallels observations in image diffusion models, where self-attention maps preserve the spatial structure of images (Liu et al. 2024a). To validate this, we conducted self-attention map replacement experiments. Fig. 2 shows that replacing self-attention maps in layers 8-14 successfully changes attributes while preserving original melody and rhythm—the edited violin music maintains the source drum’s beat pattern. Without replacement, the music loses source temporal structure and resembles direct generation from the target prompt. However, replacing all layers (1-16) partially preserves original timbre instead of complete transformation to violin. These experimental results support our idea that self-attention maps are essential for preserving temporal structure during music editing, with selective layer replacement showing promising results.
@@
Building on our exploration of attention layers, the fundamental innovation of our approach lies in preserving the temporal structure of the original music clip $x_{0}$ within self-attention map manipulation. Based on this, we propose a straightforward yet effective approach named Melodia.
As shown in Fig. 3, in our approach, we first obtain the latent $z_{0}$ of the original music clip via a VAE (Kingma, Welling et al. 2013) encoder $\mathcal{E}$, and then collect the SA features of $z_{0}$ during the Partial DDIM Inversion (Song, Meng, and Ermon 2020; Manor and Michaeli 2024) process (hereafter DDIM inversion), a version that applies the inversion only up to $T_{\rm start}\in[0,T],T=1000$ (Manor and Michaeli 2024). Specifically, for predefined time steps $t\in\{0,\dots,T_{\rm start}\}$, we invert the original latent $z_{0}$ into an editable intermediate latent $z_{T_{\rm start}}$ and store self-attention queries $Q^{\rm s}_{t}$ and keys $K^{\rm s}_{t}$ at each time step to build an attention repository in this process. The repository storing self-attention features of the original music clip is used to provide guidance for subsequent music editing. In the music editing process, we transform the features stored in the attention repository into original music clip structural guidance through the proposed Attention-based Structure Retention (ASR) method, which is based on manipulation of the self-attention mechanisms of pre-trained diffusion models, without extra fine-tuning or new blocks. Similarly to the guidance of the given target prompt, this guidance also modulates the denoising process. Finally, the output $z^{\prime}_{0}$ is decoded into a readable mel spectrogram of the desired music clip via the corresponding VAE decoder $\mathcal{D}$.
@@
Based on the above assumption, an intuitive illustration is presented in Fig. 4, showing a Structure-Timbre sampling space of the music latent diffusion model (Liu et al. 2023, 2024c). In our music editing process, the starting latent $z^{\prime}_{T_{\rm start}}$, copied from DDIM Inversion (Song, Meng, and Ermon 2020) prior $z_{T_{\rm start}}$, provides implicit structure guidance for generation. However, as shown in Fig. 4 (a), relying on this implicit guidance leads to significant structure divergence in the editing process (Song, Meng, and Ermon 2020; Mou et al. 2023; Tumanyan et al. 2023). Since the target prompt $y$ contains rich semantics of music, it can provide strong semantic guidance for editing, which is significantly stronger than the implicit structure guidance. Therefore, we build an attention repository to store SA features at each inversion time step in the system memory, and use the stored features to provide explicit structure guidance for the editing process, reducing divergence (shown in Fig. 4 (b)).
@@
where $\phi^{\prime\rm s}_{t}$ is the output of a manipulated self-attention layer and $V^{\prime\rm s}_{t}$ is the projected self-attention values of target latent $z^{\prime}_{t}$ at the specific time step $t$. In addition, we apply this manipulation to layers 8-14 of AudioLDM 2 (Liu et al. 2024c). Layer selection is analyzed in the Experiments section.
@@
MusicDelta ZoME-Bench MelodiaEdit
@@
For subjective evaluation, we introduce the Music Editing Balance (MEB) metric to capture human perception. Additionally, adapted from MusicMagus (Zhang et al. 2024b), we employ Relevance (REL) to assess how well the edited music aligns with the target prompt, and Structural Consistency (CON) to evaluate the consistency of the pitch contour and structural aspects. These metrics are rated on a five-point Likert scale (Likert 1932), with higher scores being better.
Figure 5: Quantitative Comparison with methods over $T_{\text{start}}$ range of 300-1000 on MelodiaEdit. The highlighted region is the optimal balance region where shows both text adherence and structural integrity. Our method outperforms other approaches across all $T_{\text{start}}$ values.
@@
We apply Z-score normalization $\mathcal{N}(\cdot)$ and Min-Max scaling $\mathcal{S}(\cdot)$ to ensure scores fall within $[0,1]$, with 1 indicating best performance and 0 the worst across evaluated methods.
@@
We evaluate our method using MusicDelta (Bittner et al. 2014), ZoME-Bench (Liu et al. 2024b), and MelodiaEdit, our newly constructed benchmark. MelodiaEdit comprises 2,015 music-prompt pairs designed for comprehensive intra-stem editing evaluation. The dataset contains five subsets: three synthesized subsets with 1,090 pairs generated via AudioLDM 2 (Liu et al. 2024c), and two authentic subsets with 925 pairs reconstructed from Medley-solos-DB (lostanlen2016realins) and GTZAN datasets.
@@
We evaluate our proposed method through a comparison with five competing approaches. All diffusion-based methods are implemented using the pre-trained AudioLDM 2 (Liu et al. 2024c) model with 200 inference steps. The target CFG of our method is set to 5.5.
Quantitative Comparison with baselines on datasets. Table 3 compares our method with baseline approaches. Our method achieves competitive CLAP and the lowest LPAPS across all datasets, indicating superior text adherence and temporal structure preservation. Melodia also maintains good Chroma and FAD performance, effectively preserving harmonic elements and audio quality. While MusicMagus (Zhang et al. 2024b) shows high Chroma and FAD on MelodiaEdit, this indicates editing failure as outputs remain unchanged from the source music, explaining its poor textual adherence. The ASB and AMB results demonstrate the necessity of our proposed composite metrics. Several baselines achieve 0.00 on these metrics due to severe imbalances between text adherence and structure preservation, while Melodia achieves excellent performance on both composite metrics, demonstrating superior balance between editing effectiveness and structure preservation.
@@
MusicDelta zoME-Bench MelodiaEdit
@@
Quantitative Comparison with baselines across $T_{\text{start}}$s. To quantitatively evaluate the adherence of edited music to the target prompt and the fidelity to the original music, we plot the CLAP-LPAPS results on MelodiaEdit for all methods in Fig. 5. MusicMagus (Zhang et al. 2024b) is tested only at $T_{\text{start}}=1000$ following its original setting, while other diffusion-based methods are evaluated across multiple $T_{\text{start}}$ values ranging from 300-1000. The results show that MusicGen (Copet et al. 2023) achieves good textual adherence but poor structural preservation. Notably, other diffusion-based methods also fail to precisely preserve the structure of the source music. In contrast, Melodia addresses this issue by introducing structure guidance and achieves better LPAPS scores for any level of CLAP.
@@
Qualitative Evaluation. Fig. 6 compares editing spectrograms between Melodia and baseline methods. Our method successfully preserves the temporal structure of the original music while effectively transferring the target attributes, which can be verified by more consistent rhythmic patterns maintained from the source music and successful frequency pattern transfer. All methods use the same hyperparameter settings as in the quantitative evaluation.
@@
We implemented subjective listening evaluations across three datasets. A total of 100 participants were recruited from the MIR community and music forums, with 77 valid responses retained after screening. All participants had some degree of musical experience. The test is approved by the Institutional Review Board (IRB). We randomly selected 10 test samples from each dataset, and participants rated the edited results from each method. As shown in Table 4, Melodia substantially outperforms all baseline methods on the MEB metric, indicating that our method successfully achieves the desired balance between textual alignment and structural preservation. Melodia also obtained high scores on REL and CON metrics, confirming its superior performance in semantic alignment and structure preservation.
@@
Effects of layer selection. The selection of SA replacement layers is crucial in balancing structure preservation and attribute modification. Experimental results on MelodiaEdit demonstrate that replacing SA maps in layers 8–14 of AudioLDM 2 achieves optimal performance. As shown in Tab. 5, this layer selection yields the highest balance scores (ASB and AMB) while maintaining low structural disruption (LPAPS) and high textual alignment (CLAP). Visual evidence in Fig. 2 also confirms that SA manipulation in layers 8-14 successfully preserves rhythmic patterns and melodic contours during attribute transfer.
Results on another pre-trained diffusion model. To evaluate the generalization capability of Melodia across different diffusion models, we conducted experiments with the Stable Audio Open (Evans et al. 2025) for timbre transfer on MelodiaEdit. As shown in Tab. 6, Melodia achieves improvements in both textual adherence (CLAP) and structural integrity (LPAPS and Chroma) compared to the baseline. These results confirm Melodia’s generalizability across different model architectures. Additionally, Melodia operates effectively across different sampling rates: 16kHz on AudioLDM2 (Liu et al. 2024c) and 44.1kHz on Stable Audio Open (Evans et al. 2025).